OpenStudy (anonymous):

Generate the first 5 terms of this sequence: f(1) = 0 and f(2) = 1, f(n) = f(n - 1) + f(n - 2), for n > 2. 0, -1, 1, 0, 2 0, 1, 1, 2, 3 0, 1, 2, 2, 3 0, 1, 1, 2, 2

OpenStudy (anonymous):

The next number is 3. So you plug in n = 3.

OpenStudy (anonymous):

f(3) = f(3 - 1) + f(3-2) Solve.

OpenStudy (anonymous):

That would be: f(3) = f(2) + f(1)

OpenStudy (anonymous):

Now plug in for f(2) and f(1). f(3) = 1 + 0

OpenStudy (anonymous):

f(3) = 1

OpenStudy (anonymous):

Do the same for 4 and 5. Tell me if you need help.

OpenStudy (anonymous):

Im not really good with these problems. Can you explain to me how to do it correctly?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Now we can plug in 4.

OpenStudy (anonymous):

Tell me if you need help.

OpenStudy (anonymous):

I dont mean like whats the next thing i do in this problem. I wanna know how to do any problems that are similar to this one

OpenStudy (anonymous):

Plug in 4. so, here is our original equation. f(n) = f(n - 1) + f(n - 2) n is just a variable. so for n, we plug in 4. f(4) = f(4-1) + f(4 - 2)

OpenStudy (anonymous):

Can you solve for f(4) = f(4-1) + f(4 - 2)

OpenStudy (anonymous):

f(4) = f(4) -1 + f(4) -2? Or f(3) + f(2) One of these ways?

OpenStudy (anonymous):

The second is correct!

OpenStudy (anonymous):

Now we plug in! Since we already know: f(3) = 1 f(2) = 1 f(4) = 1 + 1

OpenStudy (anonymous):

That means f(4) = 2

OpenStudy (anonymous):

Lets clean this up! f(1) = 0 f(2) = 1 f(3) = 1 f(4) = 2 f(5) = ?

OpenStudy (anonymous):

We need to find one more!

OpenStudy (anonymous):

f(4) + f(3) ?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Now, plug in the numbers.

OpenStudy (anonymous):

This last part is what gets me kind of confused. I got the f(4) + f(3) Do i go like this? f(4) = 1 f(3) =1 f(5) = 2?

OpenStudy (anonymous):

f(4) = 2 actually. We know this because when we solved for f(4) already. These problems are in a series of steps, to get one answer you have to know other so on and so on.

OpenStudy (anonymous):

See, when we solve for f(4) = f(4 - 1) + f(4 - 2) f(4) = f(3) + f(2) f(4) = 1 + 1 f(4) = 2 Do you see how we had to know what f(3) and f(2) were to know what f(4) is?

OpenStudy (anonymous):

Since we are solving for f(5) we had to know what f(4) is, and we do, because we did the equation before hand.

OpenStudy (anonymous):

so its f(5) = 3?

OpenStudy (anonymous):

Great job!

OpenStudy (anonymous):

Do you want to see what f(6) is for extra help?

OpenStudy (anonymous):

Sure. f(6) = f(6-1) + f(6 - 2) f(6) = f(5) + f(4) F(6) = 4?

OpenStudy (anonymous):

Remember that f(5) = 3 and f(4) = 2

OpenStudy (anonymous):

so, f(6) = 3 + 2 f(6) = 5

OpenStudy (anonymous):

So for f(7) f(7) = f(7-1) + f(7 - 2) f(7) = f(6) + f(5) f(7) = 5+3? Do i take the last two answers?

OpenStudy (anonymous):

Yes! Very good!

OpenStudy (anonymous):

This is known as a recursive formula.

OpenStudy (anonymous):

It needs previous answers to make the new answer.

OpenStudy (anonymous):

f(8) = f(8-1) + f(8 - 2) f(8) = f(7) + f(6) f(8) = 8 + 5?

OpenStudy (anonymous):

Great! You are getting the hang of it. Now, since it just needed the first 5, we already have the answer. Do you which one it is?

OpenStudy (anonymous):

The second option

OpenStudy (anonymous):

Great!

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

No problem! Do you want extra practice?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

f(1) = 2 and f(2) = 3, f(n) = f(1) + f(2) + f(n - 1), for n > 2. Solve for the first 5

OpenStudy (anonymous):

I am going to have to go soon, but if you need any help you can message me, :)

OpenStudy (anonymous):

Ok thanks. Ill practice these in my notebook.

OpenStudy (anonymous):

No problem. Thanks for the metal!